# A Concept-Based Explainability Framework for Large Multimodal Models

### v1

@@ Abstract
Large multimodal models (LMMs) combine unimodal encoders and large language models (LLMs) to perform multimodal tasks. Despite recent advancements towards the interpretability of these models, understanding internal representations of LMMs remains largely a mystery. In this paper, we present a novel framework for the interpretation of LMMs. We propose a dictionary learning based approach, applied to the representation of tokens. The elements of the learned dictionary correspond to our proposed concepts. We show that these concepts are well semantically grounded in both vision and text. Thus we refer to these as “multi-modal concepts”. We qualitatively and quantitatively evaluate the results of the learnt concepts. We show that the extracted multimodal concepts are useful to interpret representations of test samples. Finally, we evaluate the disentanglement between different concepts and the quality of grounding concepts visually and textually. We will publicly release our code [^1] Project webpage: https://jayneelparekh.github.io/LMM_Concept_Explainability/.
@@ 1 Introduction
For interpretability research, LMMs have largely remained unexplored. Most prior works on interpreting models that process visual data, focus on convolutional neural network (CNN) based systems and classification as the underlying task. Multimodal tasks and transformer-based architectures have both been relatively less studied. LMMs operate at the intersection of both domains. Thus, despite their rapidly growing popularity, there have been very few prior attempts at understanding representations inside an LMM [39, 33, 32].
This paper aims to bridge some of these differences and study in greater detail the intermediate representations of LMMs. To this end, motivated by the concept activation vector (CAV) based approaches for CNNs [21, 17, 15, 14], we propose a novel dictionary-learning based concept extraction method, designed for application to LMMs. Our method is used to learn a concept dictionary to understand the representations of a pretrained LMM for a given word/token of interest (Eg. ‘Dog’). For this token, we build a matrix containing the LMM’s internal representation of the token. We then linearly decompose this matrix using dictionary learning. The dictionary elements of our decomposition represent our concepts. The most interesting consequence of our method is that the learnt concepts exhibit a semantic structure that can be meaningfully grounded in both visual and textual domains. They are visually grounded by extracting the images which maximally activate these concepts. They can simultaneously be grounded in the textual domain by decoding the concept through the language model of the LMM and extracting the words/tokens they are most associated to. We refer to such concept representations as multimodal concepts. Our key contributions can be summarized as follows:
  • We propose a novel concept-based explainability framework that can be used to understand internal representations of large multimodal models. To the best of our knowledge, this is the first effort targeting multimodal models at this scale.
@@ 2 Related work
There has been an increasing interest in understanding internal representations of visual-language models (VLM) like CLIP through the lens of multimodality. Goh et al. [18] for instance discover neurons termed multimodal, that activate for certain conceptual information given images as input. Recently proposed TEXTSPAN [16] and SpLiCE [7], aim to understand representations in CLIP [36] by decomposing its visual representations on textual representations. For LMMs, Palit et al. [32] extend the causal tracing used for LLMs to analyze information across different layers in an LMM. Schwettmann et al. [39] first proposed the notion of multimodal neurons existing within the LLM part of an LMM. They term the neurons “multimodal” as they translate high-level visual information to corresponding information in text modality. The neurons are discovered by ranking them by a gradient based attribution score. Pan et al. [33] recently proposed a more refined algorithm to identify such neurons based on a different neuron importance measure that leverages architectural information of transformer MLP blocks. Instead, we propose to discover a concept structure in the token representations by learning a small dictionary of multimodally grounded concepts. Limiting the analysis to a specific token of interest allows our method to discover fine details about the token in the learnt concepts.
@@ 3.1 Background for Large Multimodal Models (LMMs)
Figure 1: Overview of our multimodal concept extraction and grounding. Given a pretrained LMM for captioning and a target token (for eg. ‘Dog’), our method extracts internal representations of $f$ about $t$, across many images. These representations are collated into a matrix $\mathbf{Z}$. We linearly decompose $\mathbf{Z}$ to learn a concept dictionary $\mathbf{U}$ and its coefficients/activations $\mathbf{V}$. Each concept $u_{k}\in\mathbf{U}$, is multimodally grounded in both visual and textual domains. For text grounding, we compute the set of most probable words $\mathbf{T}_{k}$ by decoding $u_{k}$ through the unembedding matrix $W_{U}$. Visual grounding $\mathbf{X}_{k,MAS}$ is obtained via $v_{k}$ as the set of most activating samples.
@@
We consider a general model architecture for a large multimodal model $f$, that consists of: a visual encoder $f_{V}$, a trainable connector $C$, and an LLM $f_{LM}$ consisting of $N_{L}$ layers. We assume $f$ is pretrained for captioning task with an underlying dataset $\mathcal{S}=\{(X_{i},y_{i})\}_{i=1}^{N}$ consisting of images $X_{i}\in\mathcal{X}$ and their associated caption $y_{i}\subset\mathcal{Y}$. $\mathcal{X}$ and $\mathcal{Y}$ denote the space of images and set of text tokens respectively. Note that caption $y_{i}$ can be viewed as a subset of all tokens. The input to the language model $f_{LM}$ is denoted by the sequence of tokens $h^{1},h^{2},...,h^{p}$ and the output as $\hat{y}$. The internal representation of any token at some layer $l$ and position $p$ inside $f_{LM}$ is denoted as $h^{p}_{(l)}$, with $h^{p}_{(0)}=h^{p}$. For the multimodal model, the input sequence of tokens for $f_{LM}$ consists of the concatenation of: (1) $N_{V}$ visual tokens provided by the visual encoder $f_{V}$ operating on an image $X$, followed by the connector $C$, and (2) linearly embedded textual tokens previously predicted by $f_{LM}$. For $p>N_{V}$, this can be expressed as:
@@
where $h^{1},\dots,h^{N_{V}}=C(f_{V}(X))$, and $h^{p}=\text{Emb}(\hat{y}^{p-1})$ for $p>N_{V}$. To start the prediction, $h^{N_{V}+1}$ is defined as the beginning of sentence token. The output token $\hat{y}^{p}$ is obtained by normalizing $h^{p}_{(N_{L})}$, followed by an unembedding layer that applies a matrix $W_{U}$ followed by a softmax. The predicted caption $\hat{y}$ consists of the predicted tokens $\hat{y}=\{\hat{y}^{p}\}_{p>N_{V}}$ until the end of sentence token.
Training . The model is trained with next token prediction objective, to generate text conditioned on images in an auto-regressive fashion. In this work we focus on models trained to "translate" images into text, or image captioning models. These models keep the visual encoder $f_{V}$ frozen and only train the connector $C$. Recent models also finetune the LLM to improve performance. However, we find the generalization of LLMs to multimodal inputs is an interesting phenomenon to understand, thus we focus on the setup where the LLM is kept frozen.
@@ 3.2 Method overview
Fig. 1 provides a visual summary of the whole pipeline. Given a pretrained LMM $f$ and a token of interest $t\in\mathcal{Y}$, our method consists of three key parts:
@@ 3.3 Representation extraction
Given any $X\in\mathbf{X}$, we propose to extract the token representation $h_{(L)}^{p}$ from a deep layer $L$, at the first position in the predicted caption $p>N_{V}$, such that $\hat{y}^{p}=t$. The representation $z_{j}\in\mathbb{R}^{B}$ of each sample $X_{j}\in\mathbf{X}$ is then stacked as columns of the matrix $\mathbf{Z}=[z_{1},...,z_{M}]\in\mathbb{R}^{B\times M}$. Note that the representations of text tokens in $f_{LM}$ can possess a meaningful multimodal structure as they combine information from visual token representations $h_{(l)}^{p},p\leq N_{V}$. In contrast to $a_{(l)}^{p}$ and $m_{(l)}^{p}$, that represent residual information at layer $l$, $h_{(L)}^{p}$ contains the aggregated information computed by the LMM till layer $L$, providing a holistic view of its computation across all previous layers.
@@ 3.4 Decomposing the representations
The representation matrix $\mathbf{Z}\approx\mathbf{U}\mathbf{V}$, is decomposed as product of two low-rank matrices $\mathbf{U}\in\mathbb{R}^{B\times K},\mathbf{V}\in\mathbb{R}^{K\times M}$ of rank $K<<\min(B,M)$, where $K$ denotes the number of dictionary elements. The columns of $\mathbf{U}=[u_{1},...,u_{K}]$ are the basis vectors which we refer to as concept-vectors/concepts. The rows of $\mathbf{V}$ or columns of $\mathbf{V}^{T}=[v_{1},...,v_{M}],v_{i}\in\mathbb{R}^{M}$ denote the activations of $u_{i}$ for each sample. This decomposition, as previously studied in [14], can be optimized with various constraints on $\mathbf{U},\mathbf{V}$, each leading to a different dictionary. The most common ones include PCA (constraint: $\mathbf{U}^{T}\mathbf{U}=\mathbf{I}$), K-Means (constraint: columns of $\mathbf{V}$ correspond to columns of identity matrix) and NMF (constraint: $\mathbf{Z},\mathbf{U},\mathbf{V}\geq 0$). However, for our use case, NMF is not applicable as token representations $\mathbf{Z}$ do not satisfy the non-negativity constraint. Instead, we propose to employ a relaxed version of NMF, Semi-NMF [10], which allows the decomposition matrix $\mathbf{Z}$ and basis vectors $\mathbf{U}$ to contain mixed values, and only forces the activations $\mathbf{V}$ to be non-negative. Since we expect only a small number of concepts to be present in any given sample, we also encourage sparsity in activations $\mathbf{V}$. The optimization problem to decompose $\mathbf{Z}$ via Semi-NMF can be summarized as:
@@ 4 Experiments
We experiment with the DePALM model [42] that is trained for captioning task on COCO dataset [26]. The model consists of a frozen ViT-L/14 CLIP [37] encoder as the visual encoder $f_{V}$. It is followed by a transformer connector to compress the encoding into $N_{V}=10$ visual tokens. The language model $f_{LM}$ is a frozen OPT-6.7B [46] and consists of 32 layers. For uniformity and fairness all the results in the main paper are reported with number of concepts $K=20$ and for token representations from $L=31$, the final layer before unembedding layer. For Semi-NMF, we set $\lambda=1$ throughout. We consider the 5 most activating samples in $\mathbf{X}_{k,MAS}$ for visual grounding for any $u_{k}$. For text grounding, we consider top-15 tokens for $\mathbf{T}_{k}$ before applying the filtering described in Sec 3.5.
The complete dataset consists of around 120,000 images for training, and 5000 each for validation and testing with 5 captions per image following the Karpathy split. We conduct our analysis separately for various common objects in the dataset: “Dog”, “Bus”, “Train”, “Cat”, “Bear”, “Baby”, “Car”, “Cake”. The extension to other classes remains straightforward. The precise details about number of samples for learning the dictionary, or testing, is available in Appendix A.
@@ 4.1 Evaluation setup
Concept extraction during inference: To evaluate the use of $\mathbf{U}^{*}$ in understanding any test sample $X$, we first estimate the top $r$ most activating concepts activations, $\tilde{u}(X)$ (Sec. 3.5). We then estimate the correspondence between a the image $X$ and the grounded words $\mathbf{T}_{k}$ of $\tilde{u}(X)$. This correspondence is estimated via two different metrics. The primary metric is the average CLIPScore [19] between $X$ and $\mathbf{T}_{k}$. This directly estimates correspondence between the test image embedding with the grounded words of the top concepts. The secondary metric is the average BERTScore (F1) [47] between the ground-truth captions $y$ associated with $X$ and the words $\mathbf{T}_{k}$. These metrics help validate the multimodal nature of the concept dictionaries. Their use is inspired from [39]. Details for their implementation is in Appendix A.
@@
Baselines: One set of methods for evaluation are the variants of our proposed approach where we employ different dictionary learning strategies: PCA, KMeans and Semi-NMF. For evaluating concept extraction on test data with CLIPScore/BERTScore we compare against the following baselines:
@@
We also report score using ground-truth captions (GT captions) instead of considering grounded words $\mathbf{T}_{k}$, to get the best possible correspondence score. The overlap/entanglement in concept dictionary is compared between the non-random baselines: Simple, PCA, K-Means and Semi-NMF. For evaluating the visual/text grounding we compare against the random words keeping the underlying set of MAS, $\mathbf{X}_{k,MAS}$, same for both.
@@ 4.2 Results and discussion
BS top-1 ($\uparrow$) 0.207 $\pm$ 0.07 0.307 $\pm$ 0.03 0.425 $\pm$ 0.10 0.437 $\pm$ 0.08 0.544 $\pm$ 0.1
Table 1: Test data mean CLIPScore and BERTScore for top-1 activating concept for all baselines on five tokens. CLIPScore denoted as CS, and BERTScore as BS. Statistical significance is in Appendix B.1. Higher scores are better. Best score in bold, second best is underlined.
@@
Tab. 1 reports the test top-1 CLIPScore/BERTScore for all baselines and Semi-NMF on different target tokens. Appendix B contains detailed results for other tokens as well as for the PCA and K-Means variants. We report the results for only the top-$1$ activating concept, as the KMeans and Simple baselines map a given representation to a single cluster/element. Notably, Semi-NMF generally outperforms the other baselines although the Simple baseline performs competitively. More generally, Semi-NMF, K-Means, and Simple tend to clearly outperform Rnd-Words, Noise-Imgs and PCA on these metrics, indicating that these systems project representations of test images to concepts whose associated grounded words correspond well with the visual content.
@@
Tab. 2 reports the overlap between concepts for Simple, PCA, KMeans and Semi-NMF systems. Interestingly, KMeans and Simple baseline perform significantly worse than Semi-NMF/PCA with a high overlap between grounded words, often exceeding 40%. PCA outperforms other methods with almost no overlap while Semi-NMF shows some overlap. Overall, Semi-NMF strikes the best balance among all the methods, in terms of learning a concept dictionary useful for understanding test image representations, but which also learns diverse and disentangled concepts. We thus conduct the further experiments with Semi-NMF as the underlying dictionary learning method.
@@
Fig. 4 shows visual and textual grounding of concepts extracted for token ‘dog’. For brevity, we select 8 out of 20 concepts for illustration. 2. Grounding for all concepts extracted for ‘dog’ and other tokens are in Appendix C.
@@
The concept visualizations/grounding for ‘Dog’ reveal interesting insights about the global structure of the LMM’s representation. Extracted concepts capture information about different aspects of a ‘dog’. The LMM separates representation of animal ‘Dog’ with a ‘hot dog’ (Concept 1). Specifically for ‘Dog’, Concepts (2), (3) capture information about color: ’black’, ’brown’. Concept (6) encodes information about ‘long hair’ of a dog, while concept (5) activates for ‘small/puppy-like’ dogs. Beyond concepts activating for specific characteristics of a ‘dog’, we also discover concepts describing their state of actions (Concept (4) ‘playing/running’), common scenes they can occur in (Concept (8), ’herd’), and correlated objects they can occur with (Concept (7), ‘cat and dog’). We observe such diverse nature of extracted concepts even for other tokens (Appendix C). The information about concepts can be inferred via both the visual images and the associated grounded words, highlighting their coherent multimodal grounding. Notably, compared to solely visual grounding as for CAVs for CNNs, the multimodal grounding eases the process to understand a concept.
@@
Fig. 5 illustrates the use of concept dictionaries (learnt via Semi-NMF) to understand test sample representations for tokens ‘Dog’, ‘Cat’ and ‘Bus’. For each sample we show the normalized activations of the three most activating concepts, and their respective multimodal grounding. Most activating concepts often capture meaningful and diverse features of a given sample. For instance, for first sample containing a ‘Dog’, the concepts for “long hair”, “small/tiny/puppy”, and “black/white color” all simultaneously activate. The grounding for first two concepts was also illustrated in Fig. 4. Additional visualizations for test samples are shown in Appendix C, wherein we qualitatively compare interpretations of Semi-NMF to K-Means, PCA variants and Simple baseline.
Figure 6: Mean CLIPScore between visual/text grounding $\mathbf{X}_{k,MAS},\mathbf{T}_{k}$ for all concepts (Semi=NMF), across different layers $L$. Results are for tokens ‘Dog’ and ‘Cat’.
@@
Additional qualitative experiments. A qualitative analysis for grounding of extracted concepts for different layers is available in Appendix D. Our method can be also be applied to understand the processing of visual/perceptual tokens inside the LMM which also exhibit this multimodal structure. The experiment for the same can be found in Appendix E.

### v3

@@ Abstract
Large multimodal models (LMMs) combine unimodal encoders and large language models (LLMs) to perform multimodal tasks. Despite recent advancements towards the interpretability of these models, understanding internal representations of LMMs remains largely a mystery. In this paper, we present a novel framework for the interpretation of LMMs. We propose a dictionary learning based approach, applied to the representation of tokens. The elements of the learned dictionary correspond to our proposed concepts. We show that these concepts are well semantically grounded in both vision and text. Thus we refer to these as “multi-modal concepts”. We qualitatively and quantitatively evaluate the results of the learnt concepts. We show that the extracted multimodal concepts are useful to interpret representations of test samples. Finally, we evaluate the disentanglement between different concepts and the quality of grounding concepts visually and textually. Our implementation is publicly available. [^1] Project page and code: https://jayneelparekh.github.io/LMM_Concept_Explainability/
@@ 1 Introduction
For interpretability research, LMMs have largely remained unexplored. Most prior works on interpreting models that process visual data, focus on convolutional neural network (CNN) based systems and classification as the underlying task. Multimodal tasks and transformer-based architectures have both been relatively less studied. LMMs operate at the intersection of both domains. Thus, despite their rapidly growing popularity, there have been very few prior attempts at understanding representations inside an LMM [42, 41, 35, 34].
This paper aims to bridge some of these differences and study in greater detail the intermediate representations of LMMs. To this end, motivated by the concept activation vector (CAV) based approaches for CNNs [22, 17, 15, 14], we propose a novel dictionary-learning based Concept eXplainability method designed for application to LMMs, titled CoX-LMM. Our method is used to learn a concept dictionary to understand the representations of a pretrained LMM for a given word/token of interest (Eg. ‘Dog’). For this token, we build a matrix containing the LMM’s internal representation of the token. We then linearly decompose this matrix using dictionary learning. The dictionary elements of our decomposition represent our concepts. The most interesting consequence of our method is that the learnt concepts exhibit a semantic structure that can be meaningfully grounded in both visual and textual domains. They are visually grounded by extracting the images which maximally activate these concepts. They can simultaneously be grounded in the textual domain by decoding the concept through the language model of the LMM and extracting the words/tokens they are most associated to. We refer to such concept representations as multimodal concepts. Our key contributions can be summarized as follows:
  • We propose a novel concept-based explainability framework CoX-LMM, that can be used to understand internal representations of large multimodal models. To the best of our knowledge, this is the first effort targeting multimodal models at this scale.
@@ 2 Related work
There has been an increasing interest in understanding internal representations of visual-language models (VLM) through the lens of multimodality. Shukor and Cord [42] analyse multimodal tokens and shows that despite being different, visual and perceptual tokens are implicitly aligned inside LLMs. Goh et al. [18] discover neurons termed multimodal, that activate for certain conceptual information given images as input. Recently proposed TEXTSPAN [16] and SpLiCE [7], aim to understand representations in CLIP [38] by decomposing its visual representations on textual representations. For LMMs, Palit et al. [34] extend the causal tracing used for LLMs to analyze information across different layers in an LMM. Schwettmann et al. [41] first proposed the notion of multimodal neurons existing within the LLM part of an LMM. They term the neurons “multimodal” as they translate high-level visual information to corresponding information in text modality. The neurons are discovered by ranking them by a gradient based attribution score. Pan et al. [35] proposed a more refined algorithm to identify such neurons based on a different neuron importance measure that leverages architectural information of transformer MLP blocks. Instead, we propose to discover a concept structure in the token representations by learning a small dictionary of multimodally grounded concepts. Limiting the analysis to a specific token of interest allows our method to discover fine details about the token in the learnt concepts.
@@ 3.1 Background for Large Multimodal Models (LMMs)
Figure 1: Overview of multimodal concept extraction and grounding in CoX-LMM. Given a pretrained LMM for captioning and a target token (for eg. ‘Dog’), our method extracts internal representations of $f$ about $t$, across many images. These representations are collated into a matrix $\mathbf{Z}$. We linearly decompose $\mathbf{Z}$ to learn a concept dictionary $\mathbf{U}$ and its coefficients/activations $\mathbf{V}$. Each concept $u_{k}\in\mathbf{U}$, is multimodally grounded in both visual and textual domains. For text grounding, we compute the set of most probable words $\mathbf{T}_{k}$ by decoding $u_{k}$ through the unembedding matrix $W_{U}$. Visual grounding $\mathbf{X}_{k,MAS}$ is obtained via $v_{k}$ as the set of most activating samples.
@@
We consider a general model architecture for a large multimodal model $f$, that consists of: a visual encoder $f_{V}$, a trainable connector $C$, and an LLM $f_{LM}$ consisting of $N_{L}$ layers. We assume $f$ is pretrained for captioning task with an underlying dataset $\mathcal{S}=\{(X_{i},y_{i})\}_{i=1}^{N}$ consisting of images $X_{i}\in\mathcal{X}$ and their associated caption $y_{i}\subset\mathcal{Y}$. $\mathcal{X}$ and $\mathcal{Y}$ denote the space of images and set of text tokens respectively. Note that caption $y_{i}$ can be viewed as a subset of all tokens. The input to the language model $f_{LM}$ is denoted by the sequence of tokens $h^{1},h^{2},...,h^{p}$ and the output as $\hat{y}$. The internal representation of any token at some layer $l$ and position $p$ inside $f_{LM}$ is denoted as $h^{p}_{(l)}$, with $h^{p}_{(0)}=h^{p}$. Note that $h^{p}_{(l)}$ is same as the residual stream representation in LLM transformers [13] at position $p$ and layer $l$. For the multimodal model, the input sequence of tokens for $f_{LM}$ consists of the concatenation of: (1) $N_{V}$ visual tokens provided by the visual encoder $f_{V}$ operating on an image $X$, followed by the connector $C$, and (2) linearly embedded textual tokens previously predicted by $f_{LM}$. For $p>N_{V}$, this can be expressed as:
@@
where $h^{1},\dots,h^{N_{V}}=C(f_{V}(X))$, and $h^{p}=\text{Emb}(\hat{y}^{p-1})$ for $p>N_{V}$, where Emb denotes the token embedding function. To start the prediction, $h^{N_{V}+1}$ is defined as the beginning of sentence token. The output token $\hat{y}^{p}$ is obtained by normalizing $h^{p}_{(N_{L})}$, followed by an unembedding layer that applies a matrix $W_{U}$ followed by a softmax. The predicted caption $\hat{y}$ consists of the predicted tokens $\hat{y}=\{\hat{y}^{p}\}_{p>N_{V}}$ until the end of sentence token.
Training . The model is trained with next token prediction objective, to generate text conditioned on images in an auto-regressive fashion. In this work we focus on models trained to "translate" images into text, or image captioning models. These models keep the visual encoder $f_{V}$ frozen and only train the connector $C$. Recent models also finetune the LLM to improve performance. Our approach can be applied in either case, and in our experiments we consider both type of LMMs. However, we find the generalization of LLMs to multimodal inputs is an interesting phenomenon to understand, thus we focus more on the setup where the LLM is kept frozen.
@@ 3.2 Method overview
Fig. 1 provides a visual summary of the whole CoX-LMM pipeline. Given a pretrained LMM $f$ and a token of interest $t\in\mathcal{Y}$, our method consists of three key parts:
@@ 3.3 Representation extraction
Given any $X\in\mathbf{X}$, we propose to extract the residual stream representation $h_{(L)}^{p}$ from a deep layer $L$, at the first position in the predicted caption $p>N_{V}$, such that $\hat{y}^{p}=t$. The representation $z_{j}\in\mathbb{R}^{B}$ of each sample $X_{j}\in\mathbf{X}$ is then stacked as columns of the matrix $\mathbf{Z}=[z_{1},...,z_{M}]\in\mathbb{R}^{B\times M}$. Note that the representations of text tokens in $f_{LM}$ can possess a meaningful multimodal structure as they combine information from visual token representations $h_{(l)}^{p},p\leq N_{V}$. In contrast to $a_{(l)}^{p}$ and $m_{(l)}^{p}$, that represent residual information at layer $l$, $h_{(L)}^{p}$ contains the aggregated information computed by the LMM till layer $L$, providing a holistic view of its computation across all previous layers.
@@ 3.4 Decomposing the representations
The representation matrix $\mathbf{Z}\approx\mathbf{U}\mathbf{V}$, is decomposed as product of two low-rank matrices $\mathbf{U}\in\mathbb{R}^{B\times K},\mathbf{V}\in\mathbb{R}^{K\times M}$ of rank $K<<\min(B,M)$, where $K$ denotes the number of dictionary elements. The columns of $\mathbf{U}=[u_{1},...,u_{K}]$ are the basis vectors which we refer to as concept-vectors/concepts. The rows of $\mathbf{V}$ or columns of $\mathbf{V}^{T}=[v_{1},...,v_{K}],v_{i}\in\mathbb{R}^{M}$ denote the activations of $u_{i}$ for each sample. This decomposition, as previously studied in [14], can be optimized with various constraints on $\mathbf{U},\mathbf{V}$, each leading to a different dictionary. The most common ones include PCA (constraint: $\mathbf{U}^{T}\mathbf{U}=\mathbf{I}$), K-Means (constraint: columns of $\mathbf{V}$ correspond to columns of identity matrix) and NMF (constraint: $\mathbf{U},\mathbf{V}\geq 0$). NMF is considered to yield most interpretable results [14]. However, for our use case, NMF is not useful as representation matrix $\mathbf{Z}$ is not non-negative. Instead, we propose to employ a relaxed version of NMF, Semi-NMF [10], which allows the decomposition matrix $\mathbf{Z}$ and basis vectors $\mathbf{U}$ to contain mixed values, and only forces the activations $\mathbf{V}$ to be non-negative. Note that given its relations to clustering algorithms [10], enforcing non-negative combinations of decompositions is still valued from an interpretability perspective. Since we expect only a small number of concepts to be present in any given sample, we also encourage sparsity in activations $\mathbf{V}$. The optimization problem to decompose $\mathbf{Z}$ via Semi-NMF can be summarized as:
@@ 4 Experiments
In the main paper, we focus on experiments with the DePALM model [45] that is trained for captioning task on COCO dataset [27]. The model consists of a frozen ViT-L/14 CLIP [39] encoder as the visual encoder $f_{V}$. It is followed by a transformer connector to compress the encoding into $N_{V}=10$ visual tokens. The language model $f_{LM}$ is a frozen OPT-6.7B [49] and consists of 32 layers. Additional experiments with LLaVA [28] are in Appendix A. For uniformity and fairness all the results in the main paper are reported with number of concepts $K=20$ and for token representations from $L=31$, the final layer before unembedding layer. For Semi-NMF, we set $\lambda=1$ throughout. We consider the 5 most activating samples in $\mathbf{X}_{k,MAS}$ for visual grounding for any $u_{k}$. For text grounding, we consider top-15 tokens for $\mathbf{T}_{k}$ before applying the filtering described in Sec 3.5.
The complete dataset consists of around 120,000 images for training, and 5000 each for validation and testing with 5 captions per image following the Karpathy split. We conduct our analysis separately for various common objects in the dataset: “Dog”, “Bus”, “Train”, “Cat”, “Bear”, “Baby”, “Car”, “Cake”. The extension to other classes/tokens remains straightforward and is discussed in Appendix D. The precise details about number of samples for learning the dictionary, or testing, is available in Appendix C. The implementation of our method is publicly available on GitHub  [^2] https://github.com/mshukor/xl-vlms
@@ 4.1 Evaluation setup
Concept extraction during inference: To evaluate the use of $\mathbf{U}^{*}$ in understanding any test sample $X$, we first estimate the top $r$ most activating concepts activations, $\tilde{u}(X)$ (Sec. 3.5). We then estimate the correspondence between the test image $X$ and the grounded words $\mathbf{T}_{k}$ of $u_{k}\in\tilde{u}(X)$. This correspondence is estimated via two different metrics. The primary metric is the average CLIPScore [20] between $X$ and $\mathbf{T}_{k}$. This directly estimates correspondence between the test image embedding with the grounded words of the top concepts. The secondary metric is the average BERTScore (F1) [50] between the ground-truth captions $y$ associated with $X$ and the words $\mathbf{T}_{k}$. These metrics help validate the multimodal nature of the concept dictionaries. Their use is inspired from [41]. Details for their implementation is in Appendix C.
@@
Baselines: One set of methods for evaluation are the variants of CoX-LMM where we employ different dictionary learning strategies: PCA, KMeans and Semi-NMF. For evaluating concept extraction on test data with CLIPScore/BERTScore we compare against the following baselines:
@@
We also report score using ground-truth captions (GT captions) instead of grounded words $\mathbf{T}_{k}$, to get the best possible correspondence score. The overlap/entanglement in concept dictionary is compared between the non-random baselines: Simple, PCA, K-Means and Semi-NMF. For evaluating the visual/text grounding we compare against the random words keeping the underlying set of MAS, $\mathbf{X}_{k,MAS}$, same for both.
@@ 4.2 Results and discussion
BS top-1 ($\uparrow$) 0.207 $\pm$ 0.07 0.307 $\pm$ 0.03 0.425 $\pm$ 0.10 0.437 $\pm$ 0.08 0.544 $\pm$ 0.10
Table 1: Test data mean CLIPScore and BERTScore for top-1 activating concept for all baselines on five tokens. CLIPScore denoted as CS, and BERTScore as BS. Statistical significance is in Appendix D. Our CoX-LMM framework is evaluated with Semi-NMF as underlying dictionary learning method. Higher scores are better. Best score in bold, second best is underlined.
@@
Tab. 1 reports the test top-1 CLIPScore/BERTScore for all baselines and Semi-NMF on different target tokens. Appendix D contains detailed results for other tokens as well as for the PCA and K-Means variants. We report the results for only the top-$1$ activating concept, as the KMeans and Simple baselines map a given representation to a single cluster/element. Notably, Semi-NMF generally outperforms the other baselines although the Simple baseline performs competitively. More generally, Semi-NMF, K-Means, and Simple tend to clearly outperform Rnd-Words, Noise-Imgs and PCA on these metrics, indicating that these systems project representations of test images to concepts whose associated grounded words correspond well with the visual content.
@@
Tab. 2 reports the overlap between concepts for Simple baseline and PCA, K-Means and Semi-NMF variants of CoX-LMM. Interestingly, K-Means and Simple baseline perform significantly worse than Semi-NMF/PCA with a high overlap between grounded words, often exceeding 40%. PCA outperforms other methods with almost no overlap while Semi-NMF shows some overlap. Overall, Semi-NMF strikes the best balance among all the methods, in terms of learning a concept dictionary useful for understanding test image representations, but which also learns diverse and disentangled concepts. Thus, for further CoX-LMM experiments, we consider Semi-NMF as the underlying dictionary learning method.
@@
Fig. 4 shows visual and textual grounding of concepts extracted for token ‘dog’. For brevity, we select 8 out of 20 concepts for illustration. 2. Grounding for all concepts extracted for ‘dog’ and other tokens are in Appendix E.
@@
The concept visualizations/grounding for ‘Dog’ reveal interesting insights about the global structure of the LMM’s representation. Extracted concepts capture information about different aspects of a ‘dog’. The LMM separates representation of animal ‘Dog’ with a ‘hot dog’ (Concept 1). Specifically for ‘Dog’, Concepts (2), (3) capture information about color: ’black’, ’brown’. Concept (6) encodes information about ‘long hair’ of a dog, while concept (5) activates for ‘small/puppy-like’ dogs. Beyond concepts activating for specific characteristics of a ‘dog’, we also discover concepts describing their state of actions (Concept (4) ‘playing/running’), common scenes they can occur in (Concept (8), ’herd’), and correlated objects they can occur with (Concept (7), ‘cat and dog’). We observe such diverse nature of extracted concepts even for other tokens (Appendix E). The information about concepts can be inferred via both the visual images and the associated grounded words, highlighting their coherent multimodal grounding. Notably, compared to solely visual grounding as for CAVs for CNNs, the multimodal grounding eases the process to understand a concept.
@@
Fig. 5 illustrates the use of concept dictionaries (learnt via Semi-NMF) to understand test sample representations for tokens ‘Dog’, ‘Cat’ and ‘Bus’. For each sample we show the normalized activations of the three most activating concepts, and their respective multimodal grounding. Most activating concepts often capture meaningful and diverse features of a given sample. For instance, for first sample containing a ‘Dog’, the concepts for “long hair”, “small/tiny/puppy”, and “black/white color” all simultaneously activate. The grounding for first two concepts was also illustrated in Fig. 4. Additional visualizations for test samples are shown in Appendix E, wherein we qualitatively compare interpretations of Semi-NMF to K-Means, PCA variants and Simple baseline.
Figure 6: Mean CLIPScore between visual/text grounding $\mathbf{X}_{k,MAS},\mathbf{T}_{k}$ for all concepts (Semi-NMF), across different layers $L$. Results are for tokens ‘Dog’ and ‘Cat’.
@@
Additional experiments and discussion. We conduct a preliminary study to analyze the polysemanticity/superposition in concept dictionaries in Appendix B. A qualitative analysis for grounding of extracted concepts for different layers is available in Appendix F. CoX-LMM can be also be applied to understand the processing of visual/perceptual tokens inside the LMM which also exhibit this multimodal structure. The experiment for the same can be found in Appendix G. Limitations of our method are discussed in Appendix H, and the broader societal impacts are discussed in Appendix I.